# Information Access of the Oppressed: A Problem-Posing Framework for Envisioning Emancipatory Information Access Platforms

### v1

@@ 2.1. Emancipatory IA
Emancipatory IA (Mitra, 2025a) is defined [^2] While the original definition (Mitra, 2025a) was applied in relation to information retrieval, we extend the scope to information access as a whole. as “as the study and development of information access methods that challenge all forms of human oppression, and situates its activities within broader collective emancipatory praxis”. The original formulation clarifies that emancipatory IA aspires for universal humanization and the elimination of all structural forms of oppression including colonialism, racism, patriarchy, casteism, transphobia, religious persecution, and ableism. This framing rejects the premise of technological determinism in favor of recognizing the agency and responsibilities of researchers and developers to affect positive societal transformations via IA technologies. It proposes a framework of practices, projects, and provocations for emancipatory IA. The practices—adapted from Wright’s emancipatory social science framework (Wright, 2020)—include: (i) Diagnosing and critiquingto identify the ways in which IA approaches and practices may contribute systemic harms and impede emancipatory struggles, (ii) imagining viable alternative futuresinformed by feminist, queer, decolonial, anti-racist, anti-casteist, anti-ableist, abolitionist, and other emancipatory epistemologies, and (iii) elaborating theories of changethat makes explicit the path to realizing alternative desirable futures recognizing current sociopolitical realities. The framework encourages the field to proactively seek out provocations (e.g., through cross-disciplinary engagements and practices of reflexivity) to challenge current ways of thinking and reimagine radical alternatives. Finally, the framework proposes a nonexhaustive list of projects including: safeguarding against capture, surveillance, manipulation, dehumanization, and inequitable outcomes, and promoting emancipatory pedagogy.
Our current work intersects with the emancipatory IA framework in several ways. Our work is closely relevant to the practices of imagining viable alternatives. The problem-posing framework detailed in Section 4 aims to provoke new imaginaries of IA systems as well as identify concrete projects for actualizing said visions. While the emancipatory IA framework enumerates several largely-independent projects motivated by broad concerns of emancipation, our work is predicated on the assumption that we want to explore these questions in the context of a single IA platform that may be part of a broader ecosystem. This assumption not only prompts us to envision more concrete artifacts and future directions, but also stresses the importance of considering how these different concerns and projects relate and intersect with each other in complex and significant ways. In doing so, it attempts to provide a concrete starting point for the field of IA to begin putting the theoretical framework of emancipatory IA to actual practice.
2.2. Related Initiatives
@@
The Open Search Foundation [^3] https://opensearchfoundation.org/about-open-search-foundation/ is engaged in projects that are particularly relevant to our work. This includes the OpenWebSearch.EU project [^4] https://openwebsearch.eu/the-project/ and the #FreeWebSearch Charter (Foundation, 2025). The OpenWebSearch.EU project started in response to the growing concerns of the concentration of power over IA platforms in the hands of a few large corporations bound by specific national interests. According to the project website: “The project has its origins in concerns over the imbalance of the search engine market. Despite being a backbone of our digital economy, web search is dominated and limited by a few gatekeepers like Google, Microsoft, Baidu or Yandex. Thus, information as public good, with free, unbiased and transparent access is not under public control anymore. This imbalance endangers democracy and limits the innovative potential of Europe’s research landscape and economy.” They furthermore clarify their intention to contribute towards “Europe’s digital sovereignty as well as promoting an open human-centered search engine market”.
Consequently, the #FreeWebSearch Charter enumerates ten fundamental principles and demands, namely: (i) Guarantee Transparency and Explainability, (ii) foster competition and independent infrastructure, (iii) strengthen privacy and enable search without tracking and profiling, (iv) give users, content creators and advertisers more control, (v) secure equal access and non-discrimination, (vi) support information plurality and multiple perspectives, (vii) take responsibility for environmental and social impact, (viii) safeguard integrity and protect against disinformation, (ix) build search competences and critical awareness, and (x) enforce democratic control and binding rules across borders.
@@
These Open Search projects and charters significantly overlap with calls for emancipatory IA and presents critical opportunity for collaborations across proponents of both approaches. However, there are also critical divergences in motivations and aspirations between the two perspectives. The Open Search Foundation is focused on Europe’s digital sovereignty and is centered on European values, principles, legislation, and standards. While, it represents important initiatives to build counter-structures to challenge US Big Tech’s outsized dominance over global IA platforms, it does not speak for the values and needs of the global south that represents 85% of the world’s population [^5] The global south’s population estimate is based on Verón (2023).. We must recognize that Eurocentrism is not a meaningful antidote to Amercian Exceptionalism, and no IA platform is truly open nor reflects the internet in all its diversity if it systemically erases the voices of the global south. This model of challenging US Big Tech reflects the values of multipolarism in which influence over our information ecosystems is concentrated in the hands of few powerful state and supernational actors within the international order (e.g., USA, EU, China, and Russia); at the cost of further disenfranchising other often-marginalized global populations.
@@
In contrast, emancipatory IA seeks epistemic justice for both the global south as well as for marginalized populations within powerful nation-states, and is saliently decolonial in its aspirations. These differences are not simply philosphical divergences, but have material consequences for platform design and research directions. For example, the absence of trustworthy centralized institution(s) to steer and sponsor the platform development and governance, exacerbates the need for decentralized technological and social infrastructure in emancipatory IA. Emancipatory IA also requires reflective participation of marginalized populations in platform development and governance in ways that an open search platform for EU does not. Emancipatory IA also espouses different approaches to addressing certain societal concerns compared to OpenWebSearch.EU and the #FreeWebSearch Charter. For example, the #FreeWebSearch Charter promotes transparency and diversity to tackle political bias and dominant narratives. However, that risks encoding false equivalence between oppressive and emancipatory politics which may contribute to algorithmic bothsidism (Mitra, 2025b). Emancipatory IA, in contrast, pushes for structural ways to counter systemic power differentials and uplift marginalized voices. That is, while the #FreeWebSearch vision promotes what may be considered a liberal model of democracy, emancipatory IA aligns with a more critical model of democracy (Vrijenhoek et al., 2021).
@@
Decentralized social media.
Adjacent to IA, in the social media landscape, there have been longstanding concerns about the degree that platforms’ unequal application of moderation rules, creating unfair outcomes for minoritized groups (Díaz and Hecht-Felella, 2021), and the degree to which platforms make decisions at the expense of their users (Milmo, Daniel, 2022). These concerns expanded to include explicit ideological capture after Elon Musk’s takeover of Twitter in 2022 (P. Leerssen (2025); B. Haggart, C. C. Radsch, C. Riley, N. Tusikov, H. Tworek, and S. Woolley (2022); Mastodon (2025); 33). In response, protocol-based decentralized social media like the ActivityPub-based Mastodon [^6] https://www.w3.org/TR/2018/REC-activitypub-20180123/ and ATProto-based Bluesky [^7] https://bsky.social/about/blog/10-18-2022-the-at-protocol, have emerged as alternatives that explicitly center resistance to elite and corporate capture, and individual and community agency in processes like feed curation and moderation. The novelty of the decentralized paradigm means that many issues are yet to be fully solved; in particular, two issues are particularly relevant to our discussion here: the degree to which nominally decentralized networks actually facilitate decentralized operation (Yu et al., 2025) and effectively avoid capture, and the tension between democratic control of sections of federated networks and the access to a (relatively) unified discourse on a centralized platform (Hendrix and Flowers, 2022; Fraser, 2025). However, decentralized social media remains under active development, and their protocol-based nature structurally encourages community development. The ATProto ecoysytem, for example, now features independently developed interoperable services that focus on blogging  [^8] https://leaflet.pub, https://pckt.blog/, scientific preprints  [^9] https://chive.pub/, photo sharing  [^10] https://www.flashes.blue/, and even source control  [^11] https://tangled.org/. Emancipatory IA should draw insights from these ongoing efforts and in turn contribute back new insights and sociotechnical innovations.
2.3. Participatory design
Central to our framework is the notion that communities and marginalized groups are not only essential in the design, building, and governance of emancipatory IA platforms, but that the technology cannot truly be emancipatory without their reflective participation. Existing models for collaboration with wider communities—such as participatory design (also referred to as co-design)—have been popularized particularly in the field of human-computer interaction (HCI). We make explicit how our framework diverges from existing notions of participatory design, and argue that these approaches maintain a notion of centralized power and responsibility, reinforcing the same structures that make technologies susceptible perpetual de-prioritization of marginalized communities, and to authoritarian capture.
Participatory design refers to building technologies with the involvement of relevant stakeholders—often end-users—such that the final design outcome reflects and incorporates their opinions, desires, and expertise. Participatory design is therefore inherently visionary; it exists for researchers and participants to cooperatively envision and build idealized systems. In practice, participatory design has been leveraged in the development of initiatives across multiple fields including tech and computing (Duarte et al., 2018; Sabiescu et al., 2014), social programs and community projects (Asad et al., 2017; Lab, 2018), and urban and environmental design (Baumann et al., 2017). Several of these initiatives, such as the Decidim platform, to facilitate citizen participation in the allocation of public funds by Barcelona’s City Council  (17), include the wider community in not just the initial construction but also the ongoing governance of the program.
Costanza-Chock (2020) describes design justice as a conscious redistribution of the “benefits and burdens” in design, cognizant of community-led and Indigenous traditional practices, and an investigation into how design can reinforce or challenge the “matrix of domination (white supremacy, heteropatriarchy, capitalism, ableism, settler colonialism, and other forms of structural inequality)”. The author highlights several pitfalls of participatory design to explain the continual de-prioritization of marginalized voices in design. Common shortcomings include selected participants being unrepresentative of all potential users, or resembling the researchers themselves, the general lack of diversity in ‘default’ end users, and the economic pressures to appeal to the most profitable users. Additionally, naive implementations of participatory design—“Add Diverse Stakeholders and Stir” (Delgado et al., 2021)—can lead to de-politicizing of well-intentioned participation, whereby a focus on the needs of individual stakeholders obscures the consideration of potential power structures or designing solutions for collective good  (Delgado et al., 2023). Further, seemingly inclusive and democratic enactments for participation (in a neoliberal approach) can incidentally reinforce existing power structures  (Costanza-Chock, 2020). The ways in which participants are selected and included to shape technology are therefore political, and contribute to an ongoing feedback-loop which progressively alienates marginalized groups through disproportionately worse usability and relevance. The consequences of worse utility can manifest—in the context of information access—as disproportionately higher censorship and filtering for marginalized groups  (Ungless et al., 2025; Griffin, 2025), among other forms of systematic exclusion.
While authoritarian capture implies the intentional “hijacking” of technology or media to manipulate mass opinion, the above limitations of participatory design discussed so far can lead to the unintentional skewing of social and political opinions, due to the structural suppression of presence and influence of marginalized groups. However, realized examples of participatory design still tend towards centralized power structures, since project owners maintain executive authority in deciding the parameters of participation, i.e., the selection and role of stakeholders in the design process  (Delgado et al., 2023), and ultimately have the power to block proposals made by stakeholders in favor of their own ideas  (Delgado et al., 2021). We argue that this upholding of central power is a crucial vulnerability of participatory design, that, in the context of an emancipatory IA system, would render it susceptible to authoritarian capture.
The principles of participatory design as outlined by Spinuzzi (2005) underscores a separation between the two groups of researchers and participants, with participants being elected to contribute by the researchers, given their relevance and potential insight to the problem. This two-camp distinction between researchers and participants is the most significant way in which our framework for emancipatory IA deviates from existing approaches of participatory design. While the two approaches are aligned in harnessing the ‘tacit’ perspectives of the wider community, our work does not distinguish between the designers and the users, nor does it promote a structure in which feedback is collected from one group and implemented by another. Rather, the emancipatory IA framework embeds a structure of collaboration which removes the primary onus for a particular group or individual, but ensures equal agency and influence of members of the community in building future IA technologies. Costanza-Chock (2020) emphasizes the necessity to ground design justice in Freire’s theory of popular education (or critical pedagogy). We expand on possible alternative collaborative structures, inspired by Freire’s work, in the next section.
@@ 3. Information Access of the Oppressed
While Freire’s work centers on pedagogy and learning, we apply his theories to modern information access, where information seekers become ‘learners’ via the search of digital information. We argue that information access platforms can and should serve as sites to enact emancipatory pedagogy: to encourage information seekers to co-opt and co-construct the IA platform, and in turn shape how information is experienced, as part of critically examining their realities within oppressive structures, and taking action to transform them. Our work begins with the observation that we typically relate to technology either as technologists (who envision and build technology) or as users (who seek information via the platform). Although developers and researchers may sometime use the technologies they build, most users of technology do not typically contribute directly to their envisioning or development. This dichotomy between technologists and users of technology mirrors the teacher-student relationship that Freire analyzes. By delegating the tasks of envisioning, designing, and governing information access platforms exclusively to technologists, we in turn elevate them to the position of gatekeepers of knowledge and hand them tremendous influence over our collective sensemaking of our place and relationships in this world; while the rest of the population is designated to be the objects of technological influence—their existence digitized for monitoring, modeling, and manipulation. As Freire writes:

### v2

@@ 2.1. Humanistic and emancipatory values in computing
In IA, critical information theory (Fuchs, 2009) calls for the study of how information and IA relates to the “processes of oppression, exploitation, and domination” and to conduct all research activities in “solidarity with the dominated and for the abolishment of domination”. Echoing similar sentiments, emancipatory IA (Mitra, 2025a) is defined [^3] While the original definition was applied in relation to information retrieval, we extend the scope to information access as a whole. as “the study and development of information access methods that challenge all forms of human oppression, and situates its activities within broader collective emancipatory praxis”. Emancipatory IA aspires for universal humanization and the elimination of all structural forms of oppression including colonialism, racism, patriarchy, casteism, transphobia, religious persecution, and ableism. Similar to other values-driven research, emancipatory IA rejects the premise of technological determinism in favor of recognizing the agency and responsibilities of researchers and developers to affect positive societal transformations via IA technologies. The framework proposes a nonexhaustive list of projects including: safeguarding against capture, surveillance, manipulation, dehumanization, and inequitable outcomes, and promoting emancipatory pedagogy. Our work is closely relevant to emancipatory IA’s practices of imagining viable alternatives. The problem-posing framework detailed in Section 4 aims to provoke new imaginaries of IA systems and identify projects for actualizing said visions. In doing so, it provides a concrete starting point for emancipatory IA to put its theories to practice.
Incorporating liberatory epistemologies is critical to the pursuit of social justice and emancipatory outcomes in computing. However, we should not valorize “universal methods” (Avle et al., 2017). Structures of oppression manifest differently for different marginalized communities that should be interpreted in light of each community’s historical relations. This motivates the need for community participation in the making (and unmaking) of technologies which we discuss next.
2.2. Community participation in technology production
[…]
PD originated from the Scandinavian “Cooperative Design” movement during the 1960s and 1970s, and emphasizes collaboration between technologists (or designers) and relevant stakeholders (often end-users) such that the final design outcome reflects and incorporates the opinions and desires of all stakeholders. PD is inherently visionary; it aims for technologists and users to cooperatively envision and build idealized systems, saliently making a political commitment to involve users in design-decision making that affects them. In practice, participatory design has been leveraged in the development of initiatives across multiple fields including tech and computing (Duarte et al., 2018; Sabiescu et al., 2014), social programs and community projects (Asad et al., 2017; Lab, 2018), and urban and environmental design (Baumann et al., 2017). Several of these initiatives, such as the Decidim platform, to facilitate citizen participation in the allocation of public funds by Barcelona’s City Council  (47), include the wider community in not just the initial construction but also the ongoing governance of the program.
PADR combines PD with Participatory Action Research (PAR) (Bradbury and Reason, 2007). PAR emphasizes collaboration between researchers and community members engaged in collective inquiry. In PAR, “communities of inquiry and action evolve and address questions and issues that are significant for those who participate as co-researchers” (Bradbury and Reason, 2007). PAR, in turn, draws from Action Research (AR) (Hayes, 2011; Rapoport, ; Lewin, 1946; Baskerville, 1999) that parallels the Freirean literacy movements (Campos and Anderso, 2022). By combining PAR with PD, PADR invites community participation beyond the design phase; it invokes an iterative approach in collaboration with the community for greater collective understanding of the problem and to affect desired social change.
Of the different existing participatory frameworks, metadesign is the closest to our conceptualization of Freirean design. Metadesign extends the collaboration with users “to include an ongoing process in which stakeholders become co-designers—not only at design time, but throughout the whole existence of the system” (Fischer, 2003). To facilitate this, it recommends that systems be underdesigned at design time allowing users to create appropriate solutions for themselves, and to treat ‘designing the design process’ as a critical activity. Both metadesign and PADR contrast with PD in that the latter process there is typically an end-point when the design is fixed and can no longer be changed by users (Fuad-Luke, 2013).
Costanza-Chock (2020) highlights several pitfalls of PD to explain the continual de-prioritization of marginalized voices. Common shortcomings include selected participants being unrepresentative of all potential users, or resembling the researchers themselves, and economic pressures to appeal to the most profitable users. Additionally, naive implementations of PD—“Add Diverse Stakeholders and Stir” (Delgado et al., 2021)—can lead to de-politicizing of well-intentioned participation, whereby a focus on individual stakeholder needs obscures the consideration of potential power structures or designing solutions for collective good  (Delgado et al., 2023). Further, seemingly inclusive and democratic enactments for participation can still reinforce existing power structures  (Costanza-Chock, 2020), reduce participation to consultation without affording users real power (Corbett et al., 2023), allow for participation-washing (Sloane et al., 2022), and result in co-option of community knowledge and labor (Birhane et al., 2022). Such outcomes may be best summarized by the quote [^5] This quote is from a French poster by Atelier Populaire (May 1968). V&A Museum collections accession number E.784-2003.: “je participe, tu participes, il participe, nous participons, vous participez . . . ils profitent” (in English: “I participate, you participate, he participates, we participate, you participate . . . they profit”).
Realized examples of each of these frameworks still tend towards centralized power and technolgists essentialism, since technologists maintain executive authority in deciding the parameters of participation (Delgado et al., 2023), and ultimately have the power to block proposals made by stakeholders in favor of their own ideas (Delgado et al., 2021). Even in metadesign, which aspires to empower users to be co-designers, this distinction remains, evident from the following quote from Fischer (2003).
  • The goal of making systems modifiable by users does not imply transferring the responsibility of good system design to the user. In general, “normal” users do not build tools of the quality a professional designer would since users are not concerned with the tool per se, but in doing their work.
The tension here emerges from a critical difference in the definition of “users” that is operationalized in our work in contrast to previous frameworks. While in previous framings a user is defined specifically by their lack of design know-how and system development expertise, we distinguish technologists and users based on who holds the power to change the technology and who needs to seek external approval. The “user” in our framework are marginalized communities who oftentimes hold relevant expertise that are rendered invisible by the WEIRD (Western, Educated, Industrialized, Rich, and Democratic) gaze. Examples like the Māori Data Governance Model (Kukutai et al., 2023) and the rich history of technological co-option in the Brazilian favelas (Nemer, 2022) challenge this dominative imaginary of unskilled masses of marginalized peoples who must be saved from the harms of technology by altruistic technologists. Instead, we believe our analysis will be better served by trying to identify the structural barriers preventing communities from building for themselves, such as the lack of resources and legitimate access to modify system internals. Such radical proposals for community participation in shaping technologies cannot succeed in a silo, and should begin by connecting the dots between ongoing tech workers movements for platform cooperativism (Scholz, 2015; Schneider, 2018; Scholz and Schneider, 2016; Scholz, 2017), solidarity economy initiatives (Bergeron et al., 2015; Egorov and Inshakov, 2021), and unionization drives (Mathenge, 2024; Greenhouse, 2021; Marshall, 2021). Finally, the removal of central authority creates a need for decentralized governance with a mandate to ensure emancipatory outcomes. We foreground these requirements in Section 3.
In IA, active participation of users are largely lacking in the context of proprietary platforms owned by big corporations. It is commonplace for these platforms to employ their own interpretations of what users want based on extensive surveillance and modeling of their behavioral signals—e.g., clicks (Jiang et al., 2013; Chuklin et al., 2015) and mouse movements (Diaz et al., 2013)—potentially complemented by limited user consultations. There has been recent calls to move beyond this behaviorism paradigm (Ekstrand and Willemsen, 2016) towards a culture of co-constructing the platform with users using PD methodologies (Ekstrand et al., 2025). Our work goes further, to move towards not just building with users, but enabling them to build for themselves, as visualized in Figure 1.
Figure 1. The spectrum of user participation in the different paradigms for IA technology development.
2.3. Resisting capture
Decentralized social media.
In the social media landscape, there have been longstanding concerns about how platforms’ unequal application of moderation rules creates unfair outcomes (Díaz and Hecht-Felella, 2021) and how platforms make decisions at the expense of their users (Milmo, Daniel, 2022). These concerns expanded to include explicit ideological capture after Musk’s takeover of Twitter in 2022 (P. Leerssen (2025); B. Haggart, C. C. Radsch, C. Riley, N. Tusikov, H. Tworek, and S. Woolley (2022); Mastodon (2025); 75). In response, protocol-based decentralized social media like the ActivityPub-based Mastodon [^6] https://www.w3.org/TR/2018/REC-activitypub-20180123/ and ATProto-based Bluesky [^7] https://bsky.social/about/blog/10-18-2022-the-at-protocol emerged as alternatives that explicitly center resistance to elite and corporate capture, and promote individual and community agency, e.g., in feed curation and moderation. The decentralized paradigm raises its own issues, including: (i) The degree to which nominally decentralized networks actually facilitate decentralized operation (Yu et al., 2025) and effectively avoid capture, (ii) the tension between democratic control of sections of federated networks and the access to a (relatively) unified discourse on a centralized platform (Fraser, 2025), and (iii) how decentralization of technological infrastructure alone tend to reproduce dominant structures of oppression (Hendrix and Flowers, 2022). However, decentralized social media remains under active development, and their protocol-based nature structurally encourages community development. The ATProto ecoysytem, for example, now features independently developed interoperable services that focus on blogging [^8] https://leaflet.pub, https://pckt.blog/, scientific preprints [^9] https://chive.pub/, photo sharing [^10] https://www.flashes.blue/, and even source control [^11] https://tangled.org/. Emancipatory IA should draw insights from these ongoing efforts and in turn contribute back new insights and sociotechnical innovations.
@@ 2.3. Resisting capture
The Open Search Foundation [^12] https://opensearchfoundation.org/about-open-search-foundation/ is engaged in projects that are particularly relevant to our work. This includes the OpenWebSearch.EU project [^13] https://openwebsearch.eu/the-project/ and the #FreeWebSearch Charter (Foundation, 2025). The OpenWebSearch.EU project started in response to the growing concerns of the concentration of power over IA platforms in the hands of a few large corporations bound by specific national interests. According to the project website: “Despite being a backbone of our digital economy, web search is dominated and limited by a few gatekeepers like Google, Microsoft, Baidu or Yandex. […] This imbalance endangers democracy and limits the innovative potential of Europe’s research landscape and economy.” They further clarify their intention to contribute towards Europe’s digital sovereignty.
@@
These Open Search projects and charters significantly overlap with calls for emancipatory IA and presents critical opportunity for collaborations across proponents of both approaches. However, there are also critical divergences in motivations and aspirations between the two perspectives. The Open Search Foundation is focused on Europe’s digital sovereignty and is centered on European values, principles, legislation, and standards. While, it represents important initiatives to build counter-structures to challenge US Big Tech’s outsized dominance over global IA platforms, it does not speak for the values and needs of the global south that represents 85% of the world’s population [^14] The global south’s population estimate is based on Verón (2023).. We should recognize that Eurocentrism is not a meaningful antidote to American Exceptionalism, and no IA platform is truly open nor reflects the internet in all its diversity if it systemically erases the voices of the global south. This model of challenging US Big Tech reflects the values of multipolarism in which influence over our information ecosystems is concentrated in the hands of few powerful state and supernational actors within the international order (e.g., USA, EU, China, and Russia); at the cost of further disenfranchising other often-marginalized global populations.
@@
In contrast, emancipatory IA seeks epistemic justice for both the global south as well as for marginalized populations within powerful nation-states, and is saliently decolonial in its aspirations. These differences are not simply philosphical divergences, but have material consequences for platform design and research directions. For example, the absence of trustworthy centralized institution(s) to steer platform maintenance and governance, exacerbates the need for decentralization in emancipatory IA. Emancipatory IA also requires reflective participation of marginalized communities in ways that an open search platform for EU does not. Further, emancipatory IA espouses different approaches to addressing some societal concerns compared to OpenWebSearch.EU and the #FreeWebSearch Charter. For example, the #FreeWebSearch Charter promotes transparency and diversity to tackle political bias and dominant narratives. However, that risks encoding false equivalence between oppressive and emancipatory politics which may contribute to algorithmic bothsidism (Mitra, 2025b). Emancipatory IA, in contrast, pushes for structural ways to counter systemic power differentials and uplift marginalized voices. That is, while the #FreeWebSearch vision promotes what may be considered a liberal model of democracy, emancipatory IA aligns with a more critical model of democracy (Vrijenhoek et al., 2021).
@@ 3. Information Access of the Oppressed
While Freire’s work centers on pedagogy and learning, we apply his theories to modern IA, where information seekers become ‘learners’ via the search of digital information. We argue that IA platforms are sites of liberatory struggle and should operationalize the Freirean approach by encouraging communities to co-opt and co-construct the IA platform, and in turn shape how information is experienced, as part of critically examining their realities within oppressive structures, and taking action to transform them. Specifically, we draw from Freire’s theories to: (i) Surface the technologist-user dichotomy, (ii) challenge the technologists-as-liberator frame, and (iii) introduce the notion of Freirean design and extend his problem-posing approach to IA technology development.
3.1. The technologist-user dichotomy
Our work begins with the observation that we typically relate to technology either as technologists (who envision and build technology) or as users (who seek information via the platform). Although technologists may sometime use the technologies they build, most users of technology do not typically contribute directly to their envisioning or development. This technologist-user dichotomy mirrors the teacher-student relationship that Freire analyzes. By delegating the tasks of envisioning, designing, and governing IA platforms exclusively to technologists, we in turn elevate them to the position of gatekeepers of knowledge and hand them tremendous influence over our collective sensemaking of our place and relationships in this world; while the rest of the population is designated to be the objects of technological influence—their existence digitized for monitoring, modeling, and manipulation. As Freire writes: